OpenStudy (anonymous):

factor using the u substitution 36=x^4-13x^2

OpenStudy (amistre64):

let u=x^2; then this takes the usual quadratic form

Parth (parthkohli):

\[\rm x^4 - 13x^2 - 36 = 0\]Now, \(\rm x^2 = u\).\[\rm u^2 - 13u - 36 = 0\]

OpenStudy (anonymous):

@Fall12-13, you didn't give a final answer, you gave an incorrect solving method. That will not work.

OpenStudy (anonymous):

Oh well you could just have corrected me with a proper method then..

OpenStudy (anonymous):

@Fall12-13 , no need, amistre and Parth already explained how to do it.

OpenStudy (anonymous):

Ah right, sorry about that.

OpenStudy (anonymous):

at last, a solution, i think

OpenStudy (anonymous):

\[that \ is: x=\pm\sqrt{\sqrt{13}+6}\]

OpenStudy (anonymous):

That does not look correct, @godfreysown

OpenStudy (anonymous):

Assuming, the quadratic is u^2 - 13u - 36, it is not factorable further, and the quadratic formula would need to be used to solve it. The problem required only factoring and not solving, so it's possible that the original equation was not copied correctly. If it were u^2-13u+36=0, then it would factor to (u-9)(u-4)=0 and if required to solve, x^2=9, and x^2=4, which would yield the four solutions {-3, -2, 2, 3}